OpenStudy (anonymous):

A boy has $1.35 in his bank in nickels and dimes. There are 9 more nickels than dimes. How many types of each coin are there?

OpenStudy (baldymcgee6):

let n = # of nickels let d = # of dimes It follows then, 1.35 = 0.05n +0.10d d+9 = n

OpenStudy (baldymcgee6):

Do you get that?

OpenStudy (baldymcgee6):

From these two equations: 1.35 = 0.05n +0.10d d+9 = n We can then say that 1.35 = 0.05(d+9) +0.10d

OpenStudy (anonymous):